OpenStudy (anonymous):

Can someone tell me if i got the right answer i need help ):? F(x)= -8x + 5 and G(x)=10x+2 what is g(x) - f(x) A. 2x-3 B. 2x+7 C. 18x-3 D. 18x + 7

OpenStudy (anonymous):

@moneybird

jimthompson5910 (jim_thompson5910):

g(x) - f(x) = (10x + 2) - (-8x + 5) g(x) - f(x) = 10x + 2 + 8x - 5 g(x) - f(x) = ???

OpenStudy (anonymous):

i obviously did not get the right answer cause it was not one of the choices what did you get?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what is 10x + 8x

OpenStudy (anonymous):

18

jimthompson5910 (jim_thompson5910):

not just 18

jimthompson5910 (jim_thompson5910):

but very close

OpenStudy (anonymous):

18x

jimthompson5910 (jim_thompson5910):

2 + (-5) = ???

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

so the final answer is ????

OpenStudy (anonymous):

18x-3 haha thank you but why did you change the sign to a plus sign when it was minus?

jimthompson5910 (jim_thompson5910):

because when you distribute the negative through on - (-8x + 5), you get +8x - 5

jimthompson5910 (jim_thompson5910):

the negative outside flips every sign inside the parenthesis

OpenStudy (anonymous):

oh that is very true! thank you!

jimthompson5910 (jim_thompson5910):

you're welcome